Elaborate the term total revenue and marginal revenue also calculate TR & MR in the given table




Price Output Total Revenue Marginal reve




20 1




18 2




16 3




14 4




12 5

1
Expert's answer
2022-03-15T16:51:09-0400

Answers

TR: Total revenue is the full amount a seller obtains from the total sale of their product or service.


"TR =Quantity \times Price"


MR: Marginal revenue is the increase in revenue obtained from the sale of one additional unit of product.


"MR=\dfrac{Change in Revenue}{Change in Quantity}"


"\begin {matrix}Price& Output &TR & MR\\20&1&20&20\\18&2&36&16\\16&3&48&12\\14&4&56&8\\12&5&60&4\end{matrix}"